# F-95000 Cergy-Pontoise, France

### v1

@@ Appendix B Proof of Proposition 8
    for every $Y\to B$ in $FD$ such that $YB\subseteq X$, $|{\sf tuples}(\sigma(B))|=1$,
  3. 3.
    for every $Y\to B$ in $FD$ such that $Y\subseteq X$ and $B\in sch(\Gamma)$, ${\sf tuples}(\sigma(B))\subseteq Sat(\Gamma(B))$.
@@
Proof. Assume first that $\sigma$ in $m\_Chase(T)$ satisfies the items of the proposition. To show that $x$ is ${\sf C\_ans}(Q)$, we consider a repair $R$ in ${\sf Rep}(T)$ and we show that ${\sf True}(\mathcal{R})$ contains a tuple $x\gamma$ where $\gamma$ is in $Sat(\Gamma)$.
To this end, we first notice that by Proposition 1, the first two items imply that $x$ is in ${\sf Cons}(\mathcal{T})$ and thus, by Theorem 4.1, $x$ is in ${\sf True}(\mathcal{R})$. Moreover, since $\Gamma$ is independent, item 1 implies that for every $B$ in $sch(\Gamma)$, there exits $b$ in $\sigma(B)$ such that $b$ is in $Sat(\Gamma(B))$. If $S_{1}$ is the set of all $B$ in $sch(\Gamma)$ such that there exists $Y\to B$ in $FD$ where $YB\subseteq sch(Q)$, let $S_{2}=sch(\Gamma)\setminus S_{1}$ and $\gamma_{2}$ a tuple in ${\sf tuples}(\sigma(S_{2}))$ such that for every $B$ in $S_{2}$, $\gamma_{2}.B\in Sat(\Gamma(B))$ (such a tuple exists because by item 1, $\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$). We show that $x\gamma_{2}$, which belongs to ${\sf tuples}(\sigma(XS_{2}))$, is in ${\sf True}(\mathcal{R})$. Indeed, by definition of $S_{2}$, if $Y\to B$ is in $FD$ and such that $YB\subseteq XS_{2}$, then $B\not\in S_{2}$. Thus, as $\mathbf{K}\cap S_{2}=\emptyset$ (because $\mathbf{K}\cap sch(\Gamma)=\emptyset$ and $S_{2}\subseteq sch(\Gamma)$), $YB\subseteq XS_{2}$ holds if and only if $YB\subseteq X$. Consequently $x\gamma_{2}$ is in ${\sf Cons}(\mathcal{T})$ because so is $x$, showing by Theorem 4.1 that $x\gamma_{2}$ belongs to ${\sf True}(\mathcal{R})$.
Now let $B$ be in $S_{1}$ and $Y\to B$ the corresponding dependency in $FD$ such that $Y\subseteq X$. By Proposition 6, $R$ can be defined using process (P) by which a tuple $t_{\varphi}(\sigma)$ from ${\sf tuples}(\sigma)$ is set to be in ${\sf True}(\mathcal{R})$. Since $sch(t_{\varphi}(\sigma))=sch(\sigma)$ and $sch(Q)\subseteq sch(\sigma)$, we write $t_{\varphi}(\sigma)$ as $x_{\sigma}\gamma_{\sigma}q$ where $x_{\sigma}=t_{\varphi}(\sigma).X$, $\gamma_{\sigma}=t_{\varphi}(\sigma).sch(\Gamma)$, and $q=t_{\varphi}(\sigma).(sch(\sigma)\setminus sch(Q))$. In this setting, we have $t_{\varphi}(\sigma).Y=t.Y$ (because, by Proposition 2(1), $\sigma(Y)$ contains one tuple), and thus, by the Lossless-join rule (which applies here because $R\models FD$) applied to $t_{\varphi}(\sigma)$, $x\gamma_{2}$ and $Y\to B$, we obtain that $x\gamma_{2}b$ is in ${\sf True}(\mathcal{R})$. Reproducing this reasoning for every $B$ in $S_{1}$, we obtain a tuple $x\gamma_{2}\gamma_{1}$ in $sch(Q)$ such that for every $B$ in $S_{1}$, $x\gamma_{2}\gamma_{1}.B$ is in $\sigma(B)$ and thus, by item 3, in $Sat(\Gamma(B))$. It follows that $x\gamma_{1}\gamma_{2}$ is in ${\sf True}(\mathcal{R})$ and that $\gamma_{1}\gamma_{2}$ is in $Sat(\Gamma)$. Thus $x$ is in ${\sf Ans}(Q^{[R]})$, and since this reasoning holds for every repair $R$ of $T$, we obtain that $x$ is in ${\sf C\_ans}(Q)$.
@@
Conversely, let $x$ be in ${\sf C\_ans}(Q)$. By Proposition 7, $x$ belongs to ${\sf C\_ans}^{+}(Q)$, meaning that there exists $t$ in ${\sf True}(\mathcal{T})$ such that $sch(Q)\subseteq sch(t)$, $t.X=x$, $x\in{\sf Cons}(\mathcal{T})$ and $t.sch(\Gamma)\in Sat(\Gamma)$. As for every $t\in{\sf True}(\mathcal{T})$, there exists $\sigma$ in $m\_Chase(T)$ such that $t$ is in ${\sf tuples}(\sigma(sch(t)))$, there must exist $\sigma$ in $m\_Chase(T)$ such that $sch(Q)\subseteq sch(\sigma)$, $x\in\sigma(X)$, $\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$ and for every $Y\to B$ in $FD$ such that $YB\subseteq X$, $|{\sf tuples}(\sigma(B))|=1$ (because $x$ is in ${\sf Cons}(\mathcal{T})$). In other words, there exists $\sigma$ in $m\_Chase(T)$ satisfying the first two items of the proposition.
@@
Denoting by $\Sigma(Q)$ the set of all these m-tuples, let $\sigma$ be in $\Sigma(Q)$ for which item 3 is not satisfied. Then, let $Y\to B$ in $FD$ such that $Y\subseteq X$, $B\in sch(\Gamma)$ and $b$ in ${\sf tuples}(\sigma(B))\setminus Sat(\Gamma(B))$. By item 1, ${\sf tuples}(\sigma(sch(Q)))$ contains $t$ such that $t.X=x$ and $t.B=b$. As $t$ is in ${\sf True}(\mathcal{T})$, by Proposition 4, there exists a repair $R$ such that $t$ is in ${\sf True}(\mathcal{R})$. Then, for every $x\gamma$ such that $\gamma$ is in $Sat(\Gamma)$, we have $t.Y=x.Y$ and $t.B\neq\gamma.B$ (since $\gamma.B\in Sat(\Gamma(B))$ and $t.B\not\in Sat(\Gamma(B))$). Consequently, $t$ and $x\gamma$ can not both belong to ${\sf True}(\mathcal{R})$ (because $R\models FD$). Therefore, if $\Sigma(Q)$ contains one m-tuple not satisfying item 3, then $x$ cannot belong to ${\sf C\_ans}(Q)$. We moreover notice that if one $\sigma$ in $\Sigma(Q)$ does not satisfy item 3 then for any other $\sigma^{\prime}$ in $\Sigma(Q)$, we have $\sigma^{\prime}(Y)=\sigma(Y)$ and thus $\sigma^{\prime}(B)=\sigma(B)$. It thus follows that $\sigma^{\prime}$ does not satisfy item 3 either. As a consequence, we obtain that if $x$ is in ${\sf C\_ans}(Q)$ then $m\_Chase(T)$ contains an m-tuple satisfying all three items in the proposition. The proof is therefore complete. $\Box$
Appendix C Proof of Proposition 10
@@
  • If $aggr$ is $min$, $max$ or $count$, then ${\sf C\_ans}(\mathcal{AQ})=[min\_ans,max\_ans]$ where $min\_ans$ and $max\_ans$ are returned by Algorithm 3.
  • If $aggr=sum$ and if ${\sf C\_ans}(\mathcal{AQ})=[glb,lub]$, then $min\_ans$ and $max\_ans$ as returned by Algorithm 3 satisfy that $min\_ans\leq glb$ and $max\_ans\geq lub$.
@@
  • For every aggregate function and every selection condition $\Gamma$, $ans^{*}$ as returned by Algorithm 3 is equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$.
@@
Proof. We separately consider two cases, depending on whether the test line 2 in Algorithm 3 succeeds or not. If the test fails, i.e., if $change\_min\_max$ as returned by the call of Compute_aggregate has value false, this means that $m\_Chase(T)$ contains no $\sigma$ such that $\mathbf{K}\cup sch(Q)\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$. By Corollary 1, this holds if and only if the consistent answer to $Q:~{}$select $K,X$ from $T$ where $\Gamma$ is empty. Thus, the test fails if and only if there exists a repair $R$ of $T$ for which ${\sf Ans}(Q^{[R]})=\emptyset$. In this case ${\sf C\_ans}(\mathcal{AQ})$ is NULL, which is expressed by the fact that the values of $min\_ans$, $max\_ans$ and $ans^{*}$ returned by the call of Compute_aggregate are respectively -dummy, +dummy and -dummy. Hence, in this case, Algorithm 3 provides the correct answer.
@@
Suppose now that the test line 2 succeeds, i.e., that the value of $change\_min\_max$ returned by the call of Compute_aggregate is true. The statement line 16 of Algorithm 2 shows that $m\_Chase(T)$ contains at least one $\sigma$ such that $\mathbf{K}\cup sch(Q)\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$. By Corollary 1, this holds if and only if the consistent answer to $Q:~{}$select $K,X$ from $T$ where $\Gamma$ is not empty. Thus, the test succeeds if and only if for every repair $R$ of $T$, ${\sf Ans}(Q^{[R]})\neq\emptyset$, in which case, $min\_ans$, $max\_ans$ and $ans^{*}$ are proper values, either values from $adom(M_{i})$ if $aggr\neq count$, or positive integers if $aggr=count$. In this case, it can be seen that the m-tuples $\sigma$ in $m\_Chase(T)$ that contribute to the construction of the interval are such that $\mathbf{K}\cup X\cup sch(\Gamma)\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$. This is so, because of our assumption on missing values in the fact table $F$ (every tuple with an $M_{i}$-value must be defined over all key attributes); and because any $\sigma$ not defined over some attributes in $X\cup sch(\Gamma)$ or such that ${\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)=\emptyset$ cannot contribute in the consistent answer to $\mathcal{AQ}$. This explains why, when the test line 7 of Algorithm 2 fails, no action is taken.
@@
On the other hand, the m-tuples $\sigma$ such that $\mathbf{K}\cup X\cup sch(\Gamma)\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$ are precisely those in $\Sigma(\mathcal{AQ})$, and recalling that $\Sigma^{+}(\mathcal{AQ})$ is the set of m-tuples in $\Sigma(\mathcal{AQ})$ such that ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$, we have the following for every $\sigma$ in $\Sigma(\mathcal{AQ})$:
@@
$(1)$ By Proposition 4, for every $t\in{\sf tuples}(\sigma)$, there exists a repair $R$ such that $t\in{\sf True}(\mathcal{R})$. Moreover, $t$ is the unique tuple over $sch(\sigma)$ in ${\sf True}(\mathcal{R})$ having $t.K$ as a $\mathbf{K}$-value. Notice that since $t.sch(\Gamma)$ may not be in $Sat(\Gamma)$, it is possible that $t.M_{i}$ does not contribute in the computation of the aggregate in $R$.
@@
Before showing that the values returned by Algorithm 3 are as stated in Definition 5, we mention that the aggregate operators $min$, $max$ and $sum$ are not defined when their argument is empty, which we write as $aggr(\emptyset)={\tt NULL}$. Otherwise, if $v_{1}$, $v_{2}$ and $v_{3}$ are values to which $aggr$ applies, then
@@
The first two properties show that aggregate values do not depend on the order the elementary values are considered and how they are grouped during the computation. Moreover, the third property shows that, if $aggr\neq count$, the higher the values, the higher the aggregate values. In our context, recalling that elementary values are values in $\sigma(M_{i})$ for $\sigma$ in $\Sigma(\mathcal{AQ})$, this last property shows that, when the aggregate is different than $count$, for a fixed $\sigma$, the least, respectively the highest, aggregate value is obtained by considering the least, respectively the highest, possible $M_{i}$-value. These values, respectively denoted by $min_{\sigma}$ and $max_{\sigma}$ are computed lines 8-14 of Algorithm 2.
@@
Since the second property shows that aggregates can be seen as operating over a set of values we recall the following standard properties that will be used in this proof. Let $S_{1}$ and $S_{2}$ be sets of values to which $aggr$ applies, and such that $S_{1}\subseteq S_{2}$, then: $\min(S_{1})\geq\min(S_{2})$, $\max(S_{1})\leq\max(S_{2})$, $count(S_{1})\geq count(S_{2})$, and if all values in $S_{1}$ or $S_{2}$ are positive, $sum(S_{1})\leq sum(S_{2})$. We notice that regarding the last property, if values in $S_{1}$ or $S_{2}$ can be positive or negative, no generic comparison can be stated.
@@
We now explain, for each aggregate, how the bounds of the interval of the consistent answer can be obtained, and see why Algorithm 3 returns (or not) these values. First, as for every $\sigma\in\Sigma^{+}(\mathcal{AQ})$, every repair contains exactly one tuple in ${\sf tuples}(\sigma)$, the $M_{i}$-values in $\sigma(M_{i})$ if $M_{i}\not\in sch(\Gamma)$, or in $\sigma(M_{i})\cap Sat(sch(\Gamma))$ otherwise, contribute in the computation of the aggregate. Moreover, since by monotonicity, $min_{\sigma}$ contributes in the computation of $min\_ans$ and $max_{\sigma}$ contributes in the computation of $max\_ans$, we have the following, given that $\Sigma^{+}(\mathcal{AQ})\subseteq\Sigma(\mathcal{AQ})$:
@@
  • For $aggr=min$, $min\_ans=\min\{min_{\sigma}~{}|~{}\sigma\in\Sigma(\mathcal{AQ})\}$ and $max\_ans=\min\{max_{\sigma}~{}|~{}\sigma\in\Sigma^{+}(\mathcal{AQ})\}$. These values are computed respectively in lines 23 and 17 of Algorithm 2.
@@
  • For $aggr=max$, $min\_ans=\max\{min_{\sigma}~{}|~{}\sigma\in\Sigma^{+}(\mathcal{AQ})\}$ and $max\_ans=\max\{max_{\sigma}~{}|~{}\sigma\in\Sigma(\mathcal{AQ})\}$. These values are computed respectively in lines 17 and 25 of Algorithm 2.
@@
  • For $aggr=count$, $min\_ans$, respectively $max\_ans$ is the cardinality of $\Sigma^{+}(\mathcal{AQ})$, respectively of $\Sigma(\mathcal{AQ})$. These values are computed respectively in lines 17 and 27 of Algorithm 2.

### v2

@@ Appendix E Proof of Proposition 9
    for every $Y\to B$ in $FD$ such that $YB\subseteq sch(Q)$:
    1. (a)
      if $B\in X$, then $|{\sf tuples}(\sigma(B))|=1$,
    2. (b)
      if $B\in sch(\Gamma)$, then ${\sf tuples}(\sigma(B))\subseteq Sat(\Gamma(B))$.
@@
Proof. Assume first that $\sigma$ in m_Chase(T)m\_Chase(T) satisfies the items of the proposition. To show that $x=\sigma(X)$ is in ${\sf C\_ans}(Q)$, we consider a repair $R$ in ${\sf Rep}(T)$ and we show that ${\sf True}(\mathcal{R})$ contains a tuple $x\gamma$ where $\gamma$ is in $Sat(\Gamma)$.
By Proposition 7, we can assume that $R$ has been obtained by applying process (P), which implies that ${\sf True}(\mathcal{R})$ contains a tuple $t_{\sigma}$ in ${\sf tuples}(\sigma)$. Moreover, by Proposition 1, item 1 and item 2(b) imply that $|\sigma(X)|=1$, thus $t_{\sigma}.X=x$ and $x$ is in ${\sf Cons}(\mathcal{T})$. By Theorem 4.1, it follows that $x$ is in ${\sf True}(\mathcal{R})$. Moreover, since $\Gamma$ is independent, item 1 implies that for every $B$ in $sch(\Gamma)$, there exits $b$ in $\sigma(B)$ such that $b$ is in $Sat(\Gamma(B))$. Depending on whether $B$ occurs as a right hand-side of a dependency in $FD$ or not, we have the following:
(1) If $B$ is the right hand-side of no dependency in $FD$ (which happens if $B$ is a key attribute in $\mathbb{K}$), then we have $|\sigma(B)|=1$ and thus $\sigma(B)=(b)$, implying that $t_{\sigma}.B$ is in $Sat(\Gamma(B))$.
(2) Otherwise, if $B$ is the right hand-side of a dependency in $FD$, by item 2(b), $t_{\sigma}.B$ is obviously in $Sat(\Gamma(B))$.
It follows that in any case, $t_{\sigma}.B$ is in $Sat(\Gamma(B))$, thus that $t_{\sigma}.sch(\Gamma)$ is in $Sat(\Gamma)$ since $\Gamma$ is independent. Hence, for $\gamma=t_{\sigma}.sch(\Gamma)$, $x\gamma$ is in ${\sf True}(\mathcal{R})$, showing that $x$ is in ${\sf C\_ans}(Q)$.
@@
Conversely, let $x$ be in ${\sf C\_ans}(Q)$. By Proposition 8, $x$ belongs to ${\sf C\_ans}^{+}(Q)$, meaning that there exists $t$ in ${\sf True}(\mathcal{T})$ such that $sch(Q)\subseteq sch(t)$, $t.X=x$, $x\in{\sf Cons}(\mathcal{T})$ and $t.sch(\Gamma)\in Sat(\Gamma)$. As for every $t$ in ${\sf True}(\mathcal{T})$, there exists $\sigma$ in m_Chase(T)m\_Chase(T) such that $t$ is in ${\sf tuples}(\sigma(sch(t)))$, there must exist $\sigma$ in m_Chase(T)m\_Chase(T) such that $sch(Q)\subseteq sch(\sigma)$, $x\in\sigma(X)$, $\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$ and for every $Y\to B$ in $FD$ such that $YB\subseteq X$, $|{\sf tuples}(\sigma(B))|=1$ (because $x$ is in ${\sf Cons}(\mathcal{T})$). In other words, there exists $\sigma$ in m_Chase(T)m\_Chase(T) satisfying tem 1 and item 2(a) of the proposition.
@@
Denoting by $\Sigma(Q)$ the set of all these m-tuples, let $\sigma$ be in $\Sigma(Q)$ for which item 2(b) is not satisfied. Then, let $Y\to B$ in $FD$ such that $B\in sch(\Gamma)$ and $b$ in ${\sf tuples}(\sigma(B))\setminus Sat(\Gamma(B))$. By item 1, ${\sf tuples}(\sigma(sch(Q)))$ contains $t$ such that $t.X=x$ and $t.B=b$. As $t$ is in ${\sf True}(\mathcal{T})$, by Proposition 5, there exists a repair $R$ such that $t$ is in ${\sf True}(\mathcal{R})$. Then, for every $x\gamma$ such that $\gamma$ is in $Sat(\Gamma)$, we have $t.Y=x.Y$ and $t.B\neq\gamma.B$ (since $\gamma.B\in Sat(\Gamma(B))$ and $t.B\not\in Sat(\Gamma(B))$). Consequently, $t$ and $x\gamma$ can not both belong to ${\sf True}(\mathcal{R})$ (because $R\models FD$). Therefore, if $\Sigma(Q)$ contains one m-tuple not satisfying item 2(b), then $x$ cannot belong to ${\sf C\_ans}(Q)$. We moreover notice that if one $\sigma$ in $\Sigma(Q)$ does not satisfy item 2(b) then for any other $\sigma^{\prime}$ in $\Sigma(Q)$, we have $\sigma^{\prime}(Y)=\sigma(Y)$ and thus $\sigma^{\prime}(B)=\sigma(B)$. It thus follows that $\sigma^{\prime}$ does not satisfy item 2(b) either. As a consequence, we obtain that if $x$ is in ${\sf C\_ans}(Q)$ then m_Chase(T)m\_Chase(T) contains an m-tuple satisfying all three items in the proposition. The proof is therefore complete. $\Box$
Appendix F Proof of Proposition 10
[…]
  1. 1.
    $(\forall\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ}))(\exists t\in{\sf tuples}(\sigma)\cap{\sf True}(\mathcal{R}_{1}))(t.sch(\Gamma)\not\in Sat(\Gamma))$,
  2. 2.
    $(\forall\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ}))(\exists t\in{\sf tuples}(\sigma)\cap{\sf True}(\mathcal{R}_{2}))(t.sch(\Gamma)\in Sat(\Gamma))$.
Proof. The proof relies on the fact that for every repair $R$ of $T$ and every $\sigma$ in $\Sigma(\mathcal{AQ})$, ${\sf True}(\mathcal{R})$ contains exactly one tuple $t$ such that $t\in{\sf tuples}(\sigma)$ and $t.K\in\sigma(\mathbb{K})$ (the existence is a consequence of Proposition 5, unicity follows from Proposition 2(1) because $|\sigma(\mathbb{K})|=1$, $\mathbb{K}$ is a key of $R$ and $R\models FD$). As a consequence of Proposition 7, the existence of repairs $R_{1}$ and $R_{2}$ is explicitly shown using the process (P).
Step P1. For every $A_{p}^{k}$ in $sch^{*}(D_{p})\cap sch(\sigma)$ such that $\sigma(K_{p})=k_{p}$ and $\varphi_{p}^{k}$ has not been assigned:
$(b)$ If $\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$ and ${\sf tuples}(\sigma(A_{p}^{k}))\not\subseteq Sat(\Gamma(A_{k}^{p}))$, choose an $A_{p}^{k}$-value $a^{k}_{p}$ in $\sigma(K_{p}A_{p}^{k})\setminus Sat(\Gamma(A_{k}^{p}))$ and let $\varphi_{p}^{k}(k_{p})=a_{p}^{k}$. If ${\sf tuples}(\sigma(A_{p}^{k}))\subseteq Sat(\Gamma(A_{k}^{p}))$, choose any $A_{p}^{k}$-value $a^{k}_{p}$ in $\sigma(K_{p}A_{p}^{k})$ and let $\varphi_{p}^{k}(k_{p})=a_{p}^{k}$.
Step P2. If $\mathbb{K}\subseteq sch(\sigma)$, for every $M_{i}$ in $\mathbb{M}\cap sch(\sigma)$, if $\varphi_{i}(k)$ (where $k=\sigma(\mathbb{K})$) has not been assigned, then
$(a)$ If $M_{i}\not\in sch(\Gamma)$ or if $M_{i}\in sch(\Gamma)$ and $\sigma\in\Sigma^{+}(\mathcal{AQ})$, choose an $M_{i}$-value $m$ in $\sigma(M_{i})$ and let $\varphi_{i}(k)=m$.
$(b)$ Otherwise (i.e., $M_{i}\in sch(\Gamma)$ and $\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$), choose an $M_{i}$-value $m$ in $\sigma(M_{i})\setminus Sat(\Gamma(M_{i}))$ for $\varphi_{i}(k)$.
Once all m-tuples in $\Sigma(\mathcal{AQ})$ have been considered, the two steps above are completed by considering all non-processed $K_{i}$- or $\mathbb{K}$-values as done in the generic description of process (P). Consequently, the corresponding set $\varphi$ in process (P) is properly defined.
Step P3. Build up the tuples $t_{\varphi}(\sigma)$ for every $\sigma$ in m_Chase(T)m\_Chase(T). We notice that, according to the definitions of $\Sigma(\mathcal{AQ})$ and $\Sigma^{+}(\mathcal{AQ})$, the following holds:
- if $\sigma\in\Sigma^{+}(\mathcal{AQ})$ then $t_{\varphi}(\sigma).sch(\Gamma)\in Sat(\Gamma)$,
- if $\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$ then $t_{\varphi}(\sigma).sch(\Gamma)\not\in Sat(\Gamma)$.
2. Regarding the item 2, the repair $R_{2}$ is obtained as above, by changing $(b)$ in Step P1 and Step P2 as follows: instead of choosing a value not in the corresponding set $Sat(\Gamma(A))$, choose a value that belongs to $Sat(\Gamma(A))$. We notice that, according to the definitions of $\Sigma(\mathcal{AQ})$ and $\Sigma^{+}(\mathcal{AQ})$, such a choice is always possible.
Then, as above, these two steps are completed by considering all non-processed $K_{i}$- or $\mathbb{K}$-values, and Step P3 is considered. In this case, by Proposition 7, we obtain a repair $R_{2}$ such that for every $\sigma$ in $\Sigma(\mathcal{AQ})$, $t_{\varphi}(\sigma).sch(\Gamma)$ is in $Sat(\Gamma)$. Hence $R_{2}$ satisfies item 2 in the proposition, and the proof is complete. $\Box$
Appendix G Proof of Proposition 11
@@
  • If aggraggr is minmin, maxmax or countcount, then ${\sf C\_ans}(\mathcal{AQ})=[min\_ans,max\_ans]$ where min_ansmin\_ans and max_ansmax\_ans are returned by Algorithm 4.
  • If $aggr=sum$ and if ${\sf C\_ans}(\mathcal{AQ})=[glb,lub]$, then min_ansmin\_ans and max_ansmax\_ans as returned by Algorithm 4 satisfy that $min\_ans\leq glb$ and $max\_ans\geq lub$.
@@
  • For every aggregate operator and every selection condition $\Gamma$, $ans^{*}$ as returned by Algorithm 4 is equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$.
@@
Proof. We separately consider two cases, depending on whether the test line 2 in Algorithm 4 succeeds or not. If the test fails, i.e., if change_min_maxchange\_min\_max as returned by the call of Compute_aggregate has value false, this means that m_Chase(T)m\_Chase(T) contains no $\sigma$ such that $\mathbb{K}\cup sch(Q)\subseteq sch(\sigma)$, $|{\sf tuples}(\sigma(X))|=1$ and ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$. By Corollary 1, this holds if and only if the consistent answer to Q:Q:~select $\mathbb{K},X$ from $T$ where $\Gamma$ is empty. Thus, the test fails if and only if there exists a repair $R$ of $T$ for which ${\sf Ans}(Q^{[R]})=\emptyset$. In this case ${\sf C\_ans}(\mathcal{AQ})$ is NULL, which is expressed by the fact that the values of min_ansmin\_ans, max_ansmax\_ans and $ans^{*}$ returned by the call of Compute_aggregate are respectively -dummy, +dummy and -dummy. Hence, in this case, Algorithm 4 provides the correct answer.
@@
Suppose now that the test line 2 succeeds, i.e., that the value of change_min_maxchange\_min\_max returned by the call of Compute_aggregate is true. The statement line 14 of Algorithm 3 shows that m_Chase(T)m\_Chase(T) contains at least one $\sigma$ such that $\mathbb{K}\cup sch(Q)\subseteq sch(\sigma)$, $|{\sf tuples}(\sigma(X))|=1$ and ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$. By Corollary 1, this holds if and only if the consistent answer to Q:Q:~select $\mathbb{K},X$ from $T$ where $\Gamma$ is not empty. Thus, the test succeeds if and only if for every repair $R$ of $T$, ${\sf Ans}(Q^{[R]})\neq\emptyset$, in which case, min_ansmin\_ans, max_ansmax\_ans and $ans^{*}$ are proper values, either values from $adom(M_{i})$ if $aggr\neq count$, or positive integers if $aggr=count$. In this case, it can be seen that the m-tuples $\sigma$ in m_Chase(T)m\_Chase(T) that contribute to the construction of the interval are such that $\mathbb{K}\cup X\cup sch(\Gamma)\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$. This is so, because of our assumption on missing values in the fact table $F$ (every tuple with an $M_{i}$-value must be defined over all key attributes); and because any $\sigma$ not defined over some attributes in $X\cup sch(\Gamma)$ or such that ${\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)=\emptyset$ cannot contribute in the consistent answer to $\mathcal{AQ}$. This explains why, when the test line 7 of Algorithm 3 fails, no action is taken.
@@
On the other hand, the m-tuples $\sigma$ such that $\mathbb{K}\cup X\cup sch(\Gamma)\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$ are precisely those in $\Sigma(\mathcal{AQ})$, and recalling that $\Sigma^{+}(\mathcal{AQ})$ is the set of m-tuples in $\Sigma(\mathcal{AQ})$ such that ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$, we have the following for every $\sigma$ in $\Sigma(\mathcal{AQ})$:
@@
$(1)$ By Proposition 5, for every $t\in{\sf tuples}(\sigma)$, there exists a repair $R$ such that $t\in{\sf True}(\mathcal{R})$. Moreover, $t$ is the unique tuple over $sch(\sigma)$ in ${\sf True}(\mathcal{R})$ having $t.\mathbb{K}$ as a $\mathbb{K}$-value. Notice that since $t.sch(\Gamma)$ may not be in $Sat(\Gamma)$, it is possible that $t.M_{i}$ does not contribute in the computation of the aggregate in $R$.
@@
Before showing that the values returned by Algorithm 4 are as stated in Definition 5, we mention that the aggregate operators minmin, maxmax and sumsum are not defined when their argument is empty, which we write as $aggr(\emptyset)={\tt NULL}$. Otherwise, if $v_{1}$, $v_{2}$ and $v_{3}$ are values to which aggraggr applies, then
@@
The first two properties show that aggregate values do not depend on the order the elementary values are considered and how they are grouped during the computation. Moreover, the third property shows that, if $aggr\neq count$, the higher the values, the higher the aggregate values. In our context, recalling that elementary values are values in $\sigma(M_{i})$ for $\sigma$ in $\Sigma(\mathcal{AQ})$, this last property shows that, when the aggregate is different than countcount, for a fixed $\sigma$, the least, respectively the highest, aggregate value is obtained by considering the least, respectively the highest, possible $M_{i}$-value. These values, respectively denoted by $min_{\sigma}$ and $max_{\sigma}$ are computed lines 8-12 of Algorithm 3.
@@
Since the second property shows that aggregates can be seen as operating over a set of values we recall the following standard properties that will be used in this proof. Let $S_{1}$ and $S_{2}$ be sets of values to which aggraggr applies, and such that $S_{1}\subseteq S_{2}$, then: $\min(S_{1})\geq\min(S_{2})$, $\max(S_{1})\leq\max(S_{2})$, $count(S_{1})\geq count(S_{2})$, and if all values in $S_{1}$ or $S_{2}$ are positive, $sum(S_{1})\leq sum(S_{2})$. We notice that regarding the last property, if values in $S_{1}$ or $S_{2}$ can be positive or negative, no generic comparison can be stated.
@@
We now explain, for each aggregate, how the bounds of the interval of the consistent answer can be obtained, and see why Algorithm 4 returns (or not) these values. First, as for every $\sigma\in\Sigma^{+}(\mathcal{AQ})$, every repair contains exactly one tuple in ${\sf tuples}(\sigma)$, the $M_{i}$-values in $\sigma(M_{i})$ if $M_{i}\not\in sch(\Gamma)$, or in $\sigma(M_{i})\cap Sat(sch(\Gamma))$ otherwise, contribute in the computation of the aggregate. Moreover, since by monotonicity, $min_{\sigma}$ contributes in the computation of min_ansmin\_ans and $max_{\sigma}$ contributes in the computation of max_ansmax\_ans, we have the following, given that $\Sigma^{+}(\mathcal{AQ})\subseteq\Sigma(\mathcal{AQ})$:
@@
  • For $aggr=min$, $min\_ans=\min\{min_{\sigma}~|~\sigma\in\Sigma(\mathcal{AQ})\}$ and $max\_ans=\min\{max_{\sigma}~|~\sigma\in\Sigma^{+}(\mathcal{AQ})\}$. These values are computed respectively in lines 23 and 18 of Algorithm 3.
@@
  • For $aggr=max$, $min\_ans=\max\{min_{\sigma}~|~\sigma\in\Sigma^{+}(\mathcal{AQ})\}$ and $max\_ans=\max\{max_{\sigma}~|~\sigma\in\Sigma(\mathcal{AQ})\}$. These values are computed respectively in lines 18 and 25 of Algorithm 3.
@@
  • For $aggr=count$, min_ansmin\_ans, respectively max_ansmax\_ans is the cardinality of $\Sigma^{+}(\mathcal{AQ})$, respectively of $\Sigma(\mathcal{AQ})$. These values are computed respectively in lines 18 and 27 of Algorithm 3.